the sum of 3 numbers is 125. the first number is 5 more than the 2nd. the second is 2 times the third. what are the three number?
Answer by Nsam(3)

X+Y+Z=125,REPRESENTIN 3 NUMBAZ
X=Y+5,
2Z=Y
SOVIN EQUATIONS
Y+5+Y+Z=125,
2Y+Z=120,
2(2Z)+Z=120
5Z=120,
Z=24.
2(24)=Y,
Y=48
48+24+X=125,
72+X=125,
X=53.
Deya 4 de #s r,53,48,24.